# PLAME: Leveraging Pretrained Language Models to Generate Enhanced Protein Multiple Sequence Alignments

### v1

@@ C.1 Case Study on Successful Designs
C.2 AlphaFold3 Folding Experiment
@@
C.3 Folding Enhancement on Average Proteins
@@
C.4 Further Ablation on MSA Filtering
@@
C.5 Failure Case Analysis
@@
C.6 De novo Protein Folding Enhancement
@@
From Table 7, we observed that PLAME experiences a slight decrease in pLDDT scores while simultaneously showing improvements in other metrics. The generated MSA visualizations in Figures 4 and 5 reveal that most generated sequences maintain $>70\%$ similarity to the query sequences. This phenomenon may be attributed to these test cases being highly Out-Of-Distribution (OOD) relative to the training dataset. Nevertheless, the diverse sampling strategy still effectively enhances the profile information of orphan proteins, resulting in substantial performance improvements. Furthermore, we visualized specific local regions where PLAME achieves superior alignment performance as measured by TMscore. Analysis revealed that across all augmented profiles, these high-performing local regions exhibit remarkable conservation, suggesting a strong correlation between sequence conservation patterns and structural alignment quality.
@@
Appendix D Discussion
D.1 Limitations
@@ D.2 Social Impact
Appendix E Structure Comparison Visualization
@@
Appendix F Augmented MSA Visualization
To provide an intuitive understanding of the MSAs generated by PLAME, we selected several representative cases for visualization. These cases demonstrate consistent improvements in folding accuracy compared to the MSAs provided by AF2 and cover a range of sequence lengths, including short ($<$100), medium (100-300), and long ($>$300) sequences, as well as cases under few-shot and zero-shot settings. For each visualization, the generated MSAs are highlighted with a black box. Additionally, the upper portion of each figure presents conservation information alongside the corresponding gap information. The protein information is provided in the left-top corner at each figure.
@@ Appendix F Augmented MSA Visualization
Appendix G Failure Case MSA Visualization

### v2

@@ D.1 Case Study on Successful Designs
D.2 AlphaFold3 Folding Experiment
@@
D.3 Folding Enhancement on Average Proteins
@@
D.4 Further Ablation on MSA Filtering
@@
D.5 Failure Case Analysis
@@
D.6 De novo Protein Folding Enhancement
@@
From Table 11, we observed that PLAME experiences a slight decrease in pLDDT scores while simultaneously showing improvements in other metrics. The generated MSA visualizations in Figures 4 and 5 reveal that most generated sequences maintain $70\%$ similarity to the query sequences. This phenomenon may be attributed to these test cases being highly Out-Of-Distribution (OOD) relative to the training dataset. Nevertheless, the diverse sampling strategy still effectively enhances the profile information of orphan proteins, resulting in substantial performance improvements. Furthermore, we visualized specific local regions where PLAME achieves superior alignment performance as measured by TMscore. Analysis revealed that across all augmented profiles, these high-performing local regions exhibit remarkable conservation, suggesting a strong correlation between sequence conservation patterns and structural alignment quality.
@@
Appendix E Discussion
E.1 Limitations
@@ E.2 Social Impact
Appendix F Structure Comparison Visualization
@@
Appendix G Augmented MSA Visualization
To provide an intuitive understanding of the MSAs generated by PLAME, we selected several representative cases for visualization. These cases demonstrate consistent improvements in folding accuracy compared to the MSAs provided by AF2 and cover a range of sequence lengths, including short (<<100), medium (100-300), and long (>>300) sequences, as well as cases under few-shot and zero-shot settings. For each visualization, the generated MSAs are highlighted with a black box. Additionally, the upper portion of each figure presents conservation information alongside the corresponding gap information. The protein information is provided in the left-top corner at each figure.
@@ Appendix G Augmented MSA Visualization
Appendix H Failure Case MSA Visualization